OpenStudy (anonymous):

How do I give the protons, electrons, and neutrons for an isotopes. for example: Iodine-127 Iodine-126 Iodine-128 please if you know help me not to answer it but also to understand it.

OpenStudy (anonymous):

hey so basically the proton number will always be 52

OpenStudy (anonymous):

the electron number would be the same as the proton number because the atoms are not charged, so they are not giving away any electrons... The neutron number would simply be the isotope number minus the proton number, so for iodine-127, the neutron number would be 127 - 52 = 75 neutrons x

OpenStudy (anonymous):

they will always have the same number of protons. Iodine has 53. So for Iodine-127 It's 127 - 53 (the # of Protons)= the number of neutrons. Iodone-126 is 126-53= neutrons and so forth. Carbon 12, for another example, is 12 - 6 protons = 6 neutrons.